OpenStudy (anonymous):

Simplify. Write each answer in standard form. 12. (2v^3 - v + 8) + (-v^3 + v - 3) = 14. (4h^3 + 3h + 1) - (-5h^3 + 6h - 2) =

OpenStudy (anonymous):

@Michele_Laino

OpenStudy (anonymous):

@carsonce

OpenStudy (michele_laino):

question #12 we can rewrite your expression as below: \[\begin{gathered} (2{v^3} - v + 8) + ( - {v^3} + v - 3) = \hfill \\ \hfill \\ = 2{v^3} - v + 8 - {v^3} + v - 3 = ... \hfill \\ \end{gathered} \] plese combine similar terms

OpenStudy (michele_laino):

please*

OpenStudy (anonymous):

@Michele_Laino 2v^3 - 2v + 8 - v^3 - 3 ?

OpenStudy (michele_laino):

we have: \[\begin{gathered} (2{v^3} - v + 8) + ( - {v^3} + v - 3) = \hfill \\ \hfill \\ = 2{v^3} - v + 8 - {v^3} + v - 3 = \hfill \\ \hfill \\ = {v^3}\left( {2 - 1} \right) + 8 - 3 + v\left( {1 - 1} \right) = ... \hfill \\ \end{gathered} \]

OpenStudy (michele_laino):

I have collected similar terms

OpenStudy (anonymous):

I'm confused can you talk me threw it. And is v^3 (2-1) + 8-3 + v (1-1) the answer?

OpenStudy (michele_laino):

it is not the answer, it is the expression after I have collected similar terms. Now: what is 2-1=...? 8-3=...? and 1-1=...?

OpenStudy (michele_laino):

hint: we have: \[\begin{gathered} (2{v^3} - v + 8) + ( - {v^3} + v - 3) = \hfill \\ \hfill \\ = 2{v^3} - v + 8 - {v^3} + v - 3 = \hfill \\ \hfill \\ = {v^3}\left( {2 - 1} \right) + 8 - 3 + v\left( {1 - 1} \right) = ... \hfill \\ \hfill \\ = {v^3} + 5 \hfill \\ \end{gathered} \]

OpenStudy (michele_laino):

since 2-1=1 so we have 1*v^3 1-1=0, so there is not the term proportional to v finally we have: 8-3=5

OpenStudy (michele_laino):

question 14) here we have to eliminate the parentheses, so we have to flip the sign for each term inside the second parenthesis, so we can write this: \[\begin{gathered} (4{h^3} + 3h + 1) - ( - 5{h^3} + 6h - 2) = \hfill \\ \hfill \\ = 4{h^3} + 3h + 1 + 5{h^3} - 6h + 2 = ...? \hfill \\ \end{gathered} \]

OpenStudy (anonymous):

I'm not sure how to solve questions 14 either. It doesn't make sense to me.

OpenStudy (michele_laino):

please add similar terms again

OpenStudy (michele_laino):

what is 4+5=...?

OpenStudy (anonymous):

9h^3 - 3h + 3 ? Is that correct

OpenStudy (michele_laino):

correct!

OpenStudy (michele_laino):

it is the same procedure that I have applied to the first question

OpenStudy (anonymous):

Okay can you help me with a few more?

OpenStudy (michele_laino):

I'm sorry, I'm receiving several requests for help, so please wait